OpenStudy (anonymous):

Score: A rectangle with a length of x - 4 and a width of 8 has a perimeter of 34. what is the valuee of x

Parth (parthkohli):

The perimeter of a triangle is given by \(2(length + width)\). So, the perimeter will be \(2(x - 4 + 8)\). As we already have the info that the perimeter is 34, we may equate:\[ 2(x - 4 + 8) = 34\]\[x - 4 + 8 = {34 \over 2} \]Etcetera.

OpenStudy (anonymous):

17 ?

OpenStudy (anonymous):

\(x - 4 + 8 = 17\) Now find x..

OpenStudy (anonymous):

What is the value of : -4 + 8 = ??

OpenStudy (anonymous):

@miadeex I am asking you ..

OpenStudy (anonymous):

444444444

OpenStudy (anonymous):

So: \(x + 4 = 17\) So now subtract both the sides by 4..

OpenStudy (anonymous):

13

OpenStudy (anonymous):

Well Done..